Missouri Journal of Mathematical Sciences Index

Problems Proposed

Ashbaugh, Mark, Vol. 1, No. 3, 40. Azarian, Mohammad K., Vol. 2, 80. Azarian, Mohammad K., Vol. 2, 140. Azarian, Mohammad K., Vol. 4, 27. Azarian, Mohammad K., Vol. 4, 88. Azarian, Mohammad K., Vol. 4, 145. Azarian, Mohammad K., Vol. 5, 39. Azarian, Mohammad K., Vol. 5, 90. Azarian, Mohammad K., Vol. 5, 131. Beltramo, Alvin, Vol. 4, 89. Beltramo, Alvin, Vol. 4, 145. Chen, Jin see Woods, Vol. 4. Cooper, Curtis, Vol. 0, 29. Cooper, Curtis and Robert E. Kennedy,

Vol. 0, 28. Cooper, Curtis and Robert E. Kennedy,

Vol. 0, 29. Cooper, Curtis and Robert E. Kennedy,

Vol. 1, No. 1, 30. Cooper, Curtis and Robert E. Kennedy,

Vol. 1, No. 1, 30. Cooper, Curtis and Robert E. Kennedy,

Vol. 1, No. 2, 29. Cooper, Curtis and Robert E. Kennedy,

Vol. 2, 81. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 45. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 93. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 150. Cooper, Curtis and Robert E. Kennedy,

Vol. 4, 28. Cooper, Curtis and Robert E. Kennedy,

Vol. 4, 145. Cooper, Curtis and Robert E. Kennedy,

Vol. 5, 41. Cooper, Curtis see Kennedy, Vol. 0, 28. Cooper, Curtis see Kennedy, Vol. 4, 88. Euler, Russell, Vol. 1, No. 1, 30. Euler, Russell, Vol. 1, No. 1, 31.

Euler, Russell, Vol. 2, 36. Euler, Russell, Vol. 2, 141. Euler, Russell, Vol. 3, 44. Euler, Russell, Vol. 3, 93. Euler, Russell, Vol. 3, 93. Euler, Russell, Vol. 3, 149. Euler, Russell, Vol. 4, 27. Euler, Russell, Vol. 4, 88. Euler, Russell, Vol. 5, 39. Euler, Russell, Vol. 5, 132. Ganapathy, Jayanthi, Vol. 2, 35. Ganapathy, Jayanthi, Vol. 3, 44. Ganapathy, Jayanthi, Vol. 5, 131. Hicks, Troy, Vol. 2, 36. Hicks, Troy, Vol. 3, 45. Kennedy, Robert E. and Curtis Cooper,

Vol. 0, 28. Kennedy, Robert E. and Curtis Cooper,

Vol. 4, 88. Kennedy, Robert E. see Cooper, Vol. 0,

28. Kennedy, Robert E. see Cooper, Vol. 0,

29. Kennedy, Robert E. see Cooper, Vol. 1,

No. 1, 30. Kennedy, Robert E. see Cooper, Vol. 1,

No. 1, 30. Kennedy, Robert E. see Cooper, Vol. 1,

No. 2. Kennedy, Robert E. see Cooper, Vol. 2. Kennedy, Robert E. see Cooper, Vol. 3,

45. Kennedy, Robert E. see Cooper, Vol. 3,

93. Kennedy, Robert E. see Cooper, Vol. 3,

150. Kennedy, Robert E. see Cooper, Vol. 4,

28. Kennedy, Robert E. see Cooper, Vol. 4,

145. Kennedy, Robert E. see Cooper, Vol. 5. Muthuvel, Kandasamy, Vol. 4, 28.

Volume 6, 1994 – Volume 10, 1998

Muthuvel, Kandasamy, Vol. 5, 90. Nanyes, Ollie, Vol. 5, 91. Palmer, Leonard L., Vol. 1, No. 3, 40. Rabinowitz, Stanley, Vol. 1, No. 2, 29. Rabinowitz, Stanley, Vol. 1, No. 3, 40. Rabinowitz, Stanley, Vol. 2, 35. Rabinowitz, Stanley, Vol. 2, 80. Rabinowitz, Stanley, Vol. 2, 140. Rabinowitz, Stanley, Vol. 3, 149. Rabinowitz, Stanley, Vol. 5, 40. Rabinowitz, Stanley, Vol. 5, 132. Redmond, Don, Vol. 2, 141. Shafer, Robert E., Vol. 1, No. 2, 29. Shafer, Robert E., Vol. 1, No. 2, 30. Taylor, James H., Vol. 1, No. 3, 39. Taylor, James H., Vol. 3, 94. Tinsley, Alvin, Vol. 5, 91. Wagon, Stan, Vol. 3, 150. Witt, Ana, Vol. 2, 81. Woods, Dale and Jin Chen, Vol. 4, 146. Ashbaugh, Mark, Vol. 1, No. 3, 40. Azarian, Mohammad K., Vol. 2, 80. Azarian, Mohammad K., Vol. 2, 140. Azarian, Mohammad K., Vol. 4, 27. Azarian, Mohammad K., Vol. 4, 88. Azarian, Mohammad K., Vol. 4, 145. Azarian, Mohammad K., Vol. 5, 39. Azarian, Mohammad K., Vol. 5, 90. Azarian, Mohammad K., Vol. 5, 131. Beltramo, Alvin, Vol. 4, 89. Beltramo, Alvin, Vol. 4, 145. Chen, Jin see Woods, Vol. 4. Cooper, Curtis, Vol. 0, 29. Cooper, Curtis and Robert E. Kennedy,

Vol. 0, 28. Cooper, Curtis and Robert E. Kennedy,

Vol. 0, 29. Cooper, Curtis and Robert E. Kennedy,

Vol. 1, No. 1, 30. Cooper, Curtis and Robert E. Kennedy,

Vol. 1, No. 1, 30.

Cooper, Curtis and Robert E. Kennedy,

Vol. 1, No. 2, 29. Cooper, Curtis and Robert E. Kennedy,

Vol. 2, 81. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 45. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 93. Cooper, Curtis and Robert E. Kennedy,

Vol. 3, 150. Cooper, Curtis and Robert E. Kennedy,

Vol. 4, 28. Cooper, Curtis and Robert E. Kennedy,

Vol. 4, 145. Cooper, Curtis and Robert E. Kennedy,

Vol. 5, 41. Cooper, Curtis see Kennedy, Vol. 0, 28. Cooper, Curtis see Kennedy, Vol. 4, 88. Euler, Russell, Vol. 1, No. 1, 30. Euler, Russell, Vol. 1, No. 1, 31. Euler, Russell, Vol. 2, 36. Euler, Russell, Vol. 2, 141. Euler, Russell, Vol. 3, 44. Euler, Russell, Vol. 3, 93. Euler, Russell, Vol. 3, 93. Euler, Russell, Vol. 3, 149. Euler, Russell, Vol. 4, 27. Euler, Russell, Vol. 4, 88. Euler, Russell, Vol. 5, 39. Euler, Russell, Vol. 5, 132. Ganapathy, Jayanthi, Vol. 2, 35. Ganapathy, Jayanthi, Vol. 3, 44. Ganapathy, Jayanthi, Vol. 5, 131. Hicks, Troy, Vol. 2, 36. Hicks, Troy, Vol. 3, 45. Kennedy, Robert E. and Curtis Cooper,

Vol. 0, 28. Kennedy, Robert E. and Curtis Cooper,

Vol. 4, 88. Kennedy, Robert E. see Cooper, Vol. 0,

28.